### v1

# ChainReaction! Structured Approach with Causal Chains as Intermediate Representations for Improved & Explainable Causal Video Question Answering

Abstract
Existing Causal-Why Video Question Answering (VideoQA) models often struggle with higher-order reasoning, relying on opaque, monolithic pipelines that entangle video understanding, causal inference, and answer generation. These black-box approaches offer limited interpretability and tend to depend on shallow heuristics. We propose a novel, modular framework that explicitly decouples causal reasoning from answer generation, introducing natural language causal chains as interpretable intermediate representations. Inspired by human cognitive models, these structured cause-effect sequences bridge low-level video content with high-level causal reasoning, enabling transparent and logically coherent inference. Our two-stage architecture comprises a Causal Chain Extractor (CCE) that generates causal chains from video-question pairs, and a Causal Chain-Driven Answerer (CCDA) that produces answers grounded in these chains. To address the lack of annotated reasoning traces, we introduce a scalable method for generating high-quality causal chains from existing datasets using large language models. We also propose CauCo, a new evaluation metric for causality-oriented captioning. Experiments on three large-scale benchmarks demonstrate that our approach not only outperforms state-of-the-art models, but also yields substantial gains in explainability, user trust, and generalization—positioning the CCE as a reusable causal reasoning engine across diverse domains.
1 Introduction
Understanding the motivations behind human actions is crucial for developing advanced systems capable of nuanced behavior analysis. Human actions are shaped by various factors, including personal experiences, emotions, social contexts, and cultural backgrounds. This complexity demands a comprehensive approach to uncovering the underlying causes of actions. In this context, Causal Video Question Answering (Causal-Why Video QA) requires models not just to recognize events, but to explain why they occurred—demanding higher-order reasoning that surpasses the descriptive QA paradigm (Li, Niu, and Zhang 2022; Xiao et al. 2021; Parmar et al. 2024; Foss et al. 2025).
However, existing Causal-Why Video QA models often draw conclusions from incomplete evidence or rely on shallow heuristics, such as matching action verbs or object nouns in vision-language embedding spaces (Parmar et al. 2024; Rawal et al. 2024; Wei et al. 2023). These models typically entangle video understanding, reasoning, and answer generation into a single monolithic process, which makes their reasoning opaque and error-prone. Moreover, many high-performing VLM models operate as ”black boxes,” offering limited interpretability into their decision-making processes.
Figure 1: Concept. (1) Existing Video ($\mathcal{V}$) Question ($\mathcal{Q}$) Answer ($\mathcal{A}$) approaches through the lens of structural causal models (SCMs), highlighting their monolithic and black-box nature. (2) In contrast, we propose a principled departure from this paradigm: a mediation mechanism based on natural language causal chains ($\mathcal{C}$). We factorize this SCM into two SCMs (3,4)—enabling structured video understanding, reasoning, and inference—leading to superior explainability and performance.
In this work, we make the case that reasoning and answering should be explicitly decoupled and modularized. We introduce a new paradigm in which causal reasoning and answer generation are handled by separate modules that communicate via natural language causal chains (Hanson 1955; Pearl and Mackenzie 2018)—structured sequences of cause-effect events that serve as intermediate reasoning steps (Figure 1). Causal chains provide logically coherent representations that bridge low-level video content with high-level causal understanding (Trabasso and Van Den Broek 1985). To our knowledge, ours is the first approach in Causal-Why VideoQA to explicitly use causal chains as interpretable intermediates, inspired by their role in human cognition and scientific explanation (Gopnik et al. 2004).
At the core of our method lies the integration of Structural Causal Models (SCMs) and Chain-of-Thought (CoT) reasoning, which enables a principled, structured decomposition of the VideoQA task. This synergy allows us to model causality in a robust and interpretable way, going beyond latent embeddings to explicitly capture causal semantics.
Our model consists of two stages: 1) a causal chain extractor (Figure 1(3)), and 2) a causal chain-driven answerer (Figure 1(4)). The causal chain extractor model learns to extract causal chains from video, conditioned on the causal-why questions. The causal chain-driven answerer model learns to generate answers to causal-why questions based on the extracted causal chains. However, a key challenge lies in training causal chain extractor. In particular, we do not have reasoning traces for training causal chain extractor. To tackle this challenge, we develop an approach to generate causal chains from existing VideoQA datasets efficiently.
Extensive experiments on three large scale datasets demonstrate: 1) causal chains are promising intermediate representations; 2) performance improvements across all three datasets; 3) human studies showed that causal chain-driven video QA enhances explainability and interpretability from multiple perspectives; 4) the causal chain extractor generalizes well to out-of-domain datasets, highlighting its potential as a reusable causal reasoning engine.
Our main contributions can be summarized as:
  • Proposing a structured framework for Causal-Why Video Question Answering that leverages natural language causal chains as intermediate representations to enhance reasoning and transparency.
  • Introducing a two-stage architecture—Causal Chain Extractor (CCE) and Causal Chain-Driven Answerer (CCDA)—that decouples video understanding from causal inference.
  • Presenting a novel, scalable and efficient method for generating causal chain annotations using large language models, enabling the augmentation of existing datasets.
  • Introducing CauCo score, a causality-oriented captioning metric.
  • Demonstrating through extensive experiments and human studies that the proposed approach outperforms state-of-the-art models while offering significant gains in explainability, user trust, and system debuggability.
  • Showing that the CCE generalizes well to out-of-domain datasets, enabling effective causal reasoning across diverse video domains.
2 Related Work
Chain-of-Thought (CoT)
is a prompting technique that improves large language model performance by decomposing tasks into intermediate reasoning steps. It has proven effective in complex tasks involving arithmetic, logic, and commonsense reasoning, particularly through few-shot prompting with exemplar rationales (Wei et al. 2022). Unlike most CoT-based methods that treat reasoning as latent or emergent, we model reasoning steps—causal chains—as explicit, structured, and learnable variables. This enables supervision, interpretability, and integration with modular architectures, breaking from the monolithic CoT paradigm.
Causal Reasoning and Structural Causal Models (SCMs).
Causality has long been a cornerstone of scientific reasoning, formalized in AI through SCMs and do-calculus (Pearl 2009). Recent work highlights the importance of causal representation learning for improving robustness and generalization in AI systems (Schölkopf et al. 2021). While SCMs and causal inference methods have been applied in contexts like debiasing (Tang et al. 2020) and counterfactual explanations (Narendra et al. 2018), they have rarely been extended to VideoQA. We address this gap by integrating SCMs into a modular framework where causal chains derived from video content support interpretable answer generation. This positions our approach as a novel instantiation of structured causal reasoning in a domain where temporal complexity often hinders transparency.
Causal Chains and Structured Reasoning in AI.
Causal chains are central to human explanations, offering intuitive, step-wise accounts of how and why events unfold (Parmar et al. 2024; Niu et al. 2025). In AI, structured representations like causal graphs have shown promise for improving reasoning and interpretability. Prior efforts, such as (Wang et al. 2024; Ayazoglu et al. 2013), investigate causal interactions in videos or use interventions for action recognition. However, these methods stop short of leveraging causal chains as intermediates for reasoning tasks. Our work formalizes causal chains as explicit bridges between raw video observations and high-level causal understanding, enabling a principled and interpretable approach to Causal-Why Video QA.
Video Causal Reasoning and Temporal Understanding.
A number of recent efforts explore causality in video understanding. For example, (Li et al. 2020; Yang et al. 2021) employ structural models to identify causal interactions in dynamic systems and moment retrieval. However, these do not address question answering. Some methods identify key video segments for question answering (Wei et al. 2023), while others model object interactions over space and time using generative models (Chen et al. 2023). Yet, causal reasoning remains largely implicit or limited to local interactions. Most relevant to our work is (Chen et al. 2025), who propose a dataset and task for identifying event-level causal diagrams in videos. Their model infers why a final event occurred but does not generate full causal chains or use them for causal question answering. By contrast, we extract event-level causal chains to directly power a dedicated inference module. Zang et al. (Zang et al. 2023) explore video-text causal links, while Su et al. (Su et al. 2023) generate causal questions from captions, but they do not model causal chains or explicitly decouple reasoning from answering.
Vision-Language Models and Their Limitations in Causality.
Recent Vision-Language Models (VLMs) such as VideoLLaMA (Zhang, Li, and Bing 2023), VideoChat2 (Li et al. 2024), and VILA (Lin et al. 2024) have advanced video QA and captioning by leveraging large-scale video-text pretraining. However, these models focus on recognizing and describing events, not explaining them. Their monolithic architectures lack interpretability and often rely on shallow correlations. Our approach complements these models by introducing causal chains as structured reasoning intermediates—enhancing both performance and transparency in Causal-Why Video QA.
Figure 2: Causal chain construction for SFT. (1) Human annotators of base datasets intuitively and implicitly make use of causal chains when writing correct answers. (2) We propose to recover these causal chains with the help of LLM using questions and correct gold answers. (3,4) Our robust causal chain generation and verification pipeline.
3 Causal Chain Construction for SFT
In this section, we propose a novel, scalable and efficient methodology for the construction of causal chains to be used for supervised finetuning (SFT) of our causal chain extractor (presented in later sections). Our goal is to generate causal chains that accurately reflect the reasoning behind human-written answers to video-based questions. Since no existing VideoQA dataset includes causal chain annotations to support our approach, we collect causal chains for three challenging causal video QA datasets: 1) NextQA (Xiao et al. 2021), 2) CausalVidQA (Li, Niu, and Zhang 2022), and 3) CausalChaos! (Parmar et al. 2024). We refer to these three datasets as the base (or source) datasets. We focus on Causal-Why QA in these base datasets. In base datasets, each sample is a paired triplet of {Video ($\mathcal{V}$), Question ($\mathcal{Q}$), Answer ($\mathcal{A}$)}. We have provided summaries of these three base datasets in the Appendix.
Preliminary I.
Human annotators of the base dataset watch the video ($\mathcal{V}$), then formulate a question ($\mathcal{Q}$) related to video; and then finally, meticulously write the correct, gold answer ($\mathcal{A}$) to the question. Further, the annotations are rigorously cross-annotator verified, resulting in multihuman agreement. Human annotators intuitively and implicitly make use of causal chains ($\mathcal{C}$) in the process of writing answers. This process is illustrated in Figure 2(1). As a result, the detailed QA pairs of our base datasets contain causal chains embedded within them.
Preliminary II.
Given the questions and the corresponding correct, gold answers, LLMs have been demonstrated to accurately recover the intermediate reasoning steps or causal chains. Illustrated in Figure 2(2).
Corollary I.
Based on preliminaries I and II, we propose that if provided with the questions and the corresponding human written correct, gold answers from the base datasets, then powerful oracle LLMs (e.g., GPT4o, DeepSeek, Gemini) can reliably recover the reasoning steps or causal chains.
Causal Chain Construction Methodology.
Following Corollary I, we propose a robust methodology for causal chain generation and verification in the following.
  1. 1.
    Causal Chain Generation Process: We leverage a powerful LLM (GPT4o). We refer to it as Oracle LLM. We prompt the Oracle LLM with question ($\mathcal{Q}$) and the corresponding human written correct, gold answer ($\mathcal{A}$) and ask it to generate a causal chain ($\mathcal{C}$) in natural language in a specified format. Specifically, we instruct the LLM to return causal chains in a structured format: [Event A] → [Event B] → [Event C]... [^1] Note that we do not train the LLMs to generate causal chains; rather, directly capitalize on their inherent knowledge, commonsense reasoning & synthesis capabilities to derive them. Full prompt provided in the Appendix. Illustrated in Figure 2(3).
  2. 2.
    Causal Chain Verification Process: Illustrated in Figure 2(4). We ensure that the generated causal chains are of high quality and accurately capture the reasoning steps using rigorous Quality Checks consisting of:
    1. (a)
      Programmatic Validation: Programs are developed for checking proper structure, formatting, completeness and length (constraint: $\leq$ 10 events) of the generated causal chains.
    2. (b)
      Cross-LLM Verification: Verifier LLM is provided with question ($\mathcal{Q}$), correct gold answer ($\mathcal{A}$) and the generated causal chain ($\mathcal{C}$) Generated causal chains are independently reviewed and verified by second LLM to assess the chain’s logical coherence, relevance, and consistency with $\mathcal{Q}$ and $\mathcal{A}$ pair. For this, we use a powerful LLM, but different than Oracle LLM (GPT4o) to avoid LLM circularity bias. Specifically, we use Gemini 2.5 as the verifier LLM.
    3. (c)
      Manual Verification against $\mathcal{V}$: Human verifiers are provided with full access to the corresponding video $\mathcal{V}$, the question $\mathcal{Q}$, and the correct gold answer $\mathcal{A}$ to ensure proper grounding of the causal chain in the video content. Chains passing the first two stages are verified by humans for logical coherence, relevance, and consistency with $\mathcal{Q}$ and $\mathcal{A}$ pair and grounded in video ($\mathcal{V}$). Human verifiers are computer science graduates and well-versed with the task and literature. Human verifiers are allowed to manually add missing details to the causal chains. In addition, human verifiers are trained to identify hallucinations (e.g., events not observable or implied in the video) during manual verification. Chains containing any such hallucinations are rejected and regenerated.
    If the chain fails at any stage, it is regenerated, until it passes all the checks. Finally, 1000 samples belonging to each base dataset are manually reviewed by the authors. On average, over 95% of the chains passed the author verification, ensuring high-quality annotations.
Our methodology allowed to generate consistent and reliable causal chains with respect to the video context. In total, we collected causal chains for 46024 samples across three base datasets. Further statistics provided in Appendix.
4 Approach
Theoretical Foundation and Motivation
Our work is grounded in the principles of Structural Causal Models (SCMs) (Pearl 2009) and Chain-of-Thought (CoT) (Wei et al. 2022) reasoning. We decompose the VideoQA task into two explicit stages: 1) causal chain extraction; and 2) causal chain-driven answering. Please refer to Figure 1. From an SCM perspective, we introduce causal chains as structured intermediate representations. While many existing VideoQA approaches can be abstracted as an undifferentiated model where video ($\mathcal{V}$) and question ($\mathcal{Q}$) jointly influence the answer ($\mathcal{A}$) as ($\mathcal{V}\rightarrow\mathcal{A}\leftarrow\mathcal{Q}$), we propose a structured mediation via a causal chain variable ($\mathcal{C}$), yielding the following SCM: $\mathcal{V}\rightarrow\mathcal{C}\rightarrow\mathcal{A},\mathcal{Q}\rightarrow\mathcal{C}\rightarrow\mathcal{A}$. Then, we factorize this SCM into: $\mathcal{V}\rightarrow\mathcal{C}\leftarrow\mathcal{Q}$ (this becomes our Causal Chain Extractor—Figure 1(3)) and $\mathcal{Q}\rightarrow\mathcal{A}\leftarrow\mathcal{C}$ (this becomes our Causal Chain-driven Answerer—Figure 1(4)). Causal Chain Extractor module would first extract $\mathcal{C}$, and pass it to the Causal Chain-Driven Answerer module. This structured factorization aligns with CoT principles by making reasoning steps explicit, and reflects SCM modularity by separating causal understanding from answer generation. As a result, our approach enables:
  • Focused processing: Unlike typical VideoQA approaches, the proposed structured approach allows more focused processing at each stage, with the outputs of each stage explicitly made available to the subsequent one. By handling smaller, more focused chunks, there is a lesser risk of overlooking important reasoning steps or drawing incorrect conclusions unlike existing approaches. This approach draws inspiration from the Chain-of-Thought philosophy, which we adapt into a structured model.
  • Improved video understanding: Modeling cause-and-effect relationships explicitly in causal chains is a dense prediction task. With detailed supervision, vision models learn to better able to capture rich video content. In contrast, typical VideoQA models are trained with a single label (the correct answer), which may be insufficient to fully characterize the video’s complexity (Bi et al. 2025).
  • Enhanced explainability: By generating human-readable causal chains as intermediate outputs, our model improves both explainability & interpretability, making the reasoning process more transparent.
Model
Thus far, we saw that our approach factorizes VideoQA problem into subproblems: causal chain extraction and causal chain-driven answerer. Now, we explain these individual modules and their training/inference in the following.
Causal Chain Extractor (CCE)
The objective of CCE module is to explicitly capture the cause-and-effect relationships or the reasoning steps from videos, conditioned on the questions, and express them in a detailed yet concise form. Generating causal chains from videos conditioned on questions constitutes a complex reasoning task. Towards that end, we propose to leverage the representational capacity of foundation models. However, generating causal chains is non-trivial task that extends beyond the native capabilities of existing vision-language foundation models. Although foundation models exhibit strong performance across a wide range of vision-language benchmarks, they demonstrate limited proficiency in causal chain generation, primarily due to the absence of structured causal supervision in their pretraining corpora. To address this limitation, a novel dataset comprising video–question–causal chain triplets is introduced, as described in Section 3. This dataset enables the supervised fine-tuning (SFT) of pretrained vision-language foundation models to align their generative capacity with the demands of structured causal inference. Through this process, the model is explicitly trained to capture event-level dependencies and produce human-interpretable causal chains grounded in both visual and linguistic context. Formally, given a video $\mathcal{V}$ and a corresponding question $\mathcal{Q}$, the model learns a mapping: $f_{CCE}:(\mathcal{V},\mathcal{Q})\rightarrow\mathcal{C}$; where $\mathcal{C}$ denotes the generated causal chain and $f_{CCE}$ represents the fine-tuned vision-language foundation model.
Figure 3: Training of both modules in stages.
Causal Chain-Driven Answerer (CCDA)
The objective of CCDA module is to select the correct answer from candidate answers, based on the questions and the causal chains. Existing large language models have shown strong performance on a variety of language-based tasks. We propose leveraging language models to implement the Causal Chain-Driven Answerer (CCDA). However, processing causal chains using language models is non-trivial and mostly not covered in their training suites, largely due to lack of explicit causal chains or step-by-step annotations of causal reasoning. Towards that end, we resort to supervised finetuning of CCDA. Specifically, CCDA receives the extracted causal chain, question, and candidate answer options (typically four to five in existing datasets) and is prompted to select the correct answer. Notably, the Causal Chain Extractor (CCE) remains frozen at this stage. Formally, given a question $\mathcal{Q}$, the corresponding causal chain $\mathcal{C}$, and answer options $\mathcal{O}$ the model learns a mapping: $f_{CCDA}:(\mathcal{Q},\mathcal{C},\mathcal{O})\rightarrow A$; where $\mathcal{A}$ denotes the selected answer option and $f_{CCDA}$ represents a finetuned language model.
Inference
During inference, the CCE predicts a causal chain from the video and question without using ground-truth chains. The CCDA then uses this chain, the question, and answer options to select the correct answer.
Causality-aware Captioning Metric
To evaluate the quality of generated causal chains, we also assess how respectful the chains are in terms of modeling causality. Since regular captioning metrics do not capture this aspect, we introduce a novel Causal Coherence Score—CauCo score, as an indicator of whether causal chains are causally coherent, i.e., events in the chains are logically connected by cause and effect. To achieve this, following a similar approach of LLM as verifiers, we SFT a LLM to identify if a chain is causally coherent. The final expected output of this evaluator is expected to be ”True” or ”False”. For SFT, we construct a set consisting of positive and negative samples. The positive samples are correct causal chains. To generate negative samples, we perturb the correct chains with a possibility of six distinct strategies, namely: actor swapping, event negation, event removal, event order reversal, semantic modification, and chain shuffling.
5 Experiments
This section presents experiments validating our hypotheses. We first establish an upper bound for causal VideoQA using our method, then compare it to state-of-the-art approaches, and finally assess explainability through human studies. The experimental setup is outlined below.
Implementation details. We use PyTorch (Paszke et al. 2019) to implement models. Noting the strong performance of VILA 1.5 and LLaMA on vision and language tasks, we adopt them as representative models in implementing our CCE and CCDA, respectively. Note that our approach is not designed for or limited to any models; practitioners may use models of their choice such as (Zhang, Li, and Bing 2023; Li et al. 2024; Lin et al. 2024; Chen et al. 2024). Our CCE is based on VILA-3B (Lin et al. 2024); and our CCDA is LLaMA-3.1-8B (Grattafiori et al. 2024) version. Further implementation details are provided in Appendix. We will release the codebase along with causal chains.
Datasets. Unless specified otherwise, we conduct experiments on the datasets as discussed in Section 3.
Tasks. We evaluate on a multiple-choice QA task with five answer options per question, only one of which is correct.
Performance metric. Following prior work (Xiao et al. 2021; Li, Niu, and Zhang 2022; Parmar et al. 2024), we use Accuracy as the performance metric. For causal chain quality, we use our CauCo score and captioning metrics: BLEU (Papineni et al. 2002), Meteor (Banerjee and Lavie 2005), ROUGE (Lin 2004), SPICE (Anderson et al. 2016) scores.
Experimental Upper bound on Performance
NextQA CVQA CausalChaos! Avg
99.70 99.85 98.65 99.40
Table 1: Experimental results (Accuracy in %) for answering based on ground truth causal chains.
Figure 4: Sensitivity to causal chains. (1) plot of performance drop vs links removed. (2) Qualitative example. OCC: original causal chains, MCC: masked chains, SA: selected answer. (3) Qualitative examples of SA changes based on hypothetical chains (HCC). Please zoom-in.
We first investigate whether causal chains can serve as effective intermediate representations. To this end, we use groundtruth causal chains instead of predicted ones, training & testing the Causal Chain-Driven Answerer on these annotations. This setting simulates an ideal scenario where causal chain generation is perfectly accurate. Thus, this experiment aims to determine the model’s performance upper bound when provided with flawless causal chains.
The results, summarized in Table 1, show near-perfect accuracy, even surpassing human performance. This finding suggests that using causal chains as intermediate representations is a promising paradigm worth further exploration.
Probing sensitivity to causal chains.
We investigated how the CCDA depended on causal chains; and how sensitive it is to the quality of causal chains. For this, we incrementally mask-out links of the causal chains. Quantitative and qualitative results in Figure 4(1,2) show how the performance degrades as masking is increased. We also tested how hypothetical causal chains influence CCDA’s answer selection in an intuitive way (Figure 4(3)).
Performance Comparison with SOTA
Baselines.
Following prior work (Xiao et al. 2021; Li, Niu, and Zhang 2022; Parmar et al. 2024), we compare against a wide range of models: 1) traditional approaches {BlindQA, EVQA, CoMem, HME, HCRN, HGA}; 2) causal approaches {MCR+HCRN, VCSR}; and 3) SOTA VLMs/Multimodal foundation approaches {MIST, BLIP-2, Video-LLaMA, VideoChat2, GPT-4o, VILA, DeepSeekVL2, Qwen-VL2.5} which excel on vision-language tasks.
Model NXQA CVQA CChQA Avg WtAvg
Traditional VideoQA approaches
BlindQA (Antol et al. 2015) 28.38 59.46 13.07 33.64 44.87
EVQA (Antol et al. 2015) 42.31 60.95 13.48 38.91 50.62
CoMem (Gao et al. 2018) 46.15 62.79 13.88 40.94 53.05
HME (Fan et al. 2019) 46.52 61.45 14.02 40.66 52.43
HCRN (Le et al. 2020) 47.00 61.61 17.00 41.86 52.93
HGA (Jiang and Han 2020) 47.00 63.51 15.36 42.00 53.88
Semi-traditional approaches
MIST (Gao et al. 2023) 54.79 72.41 44.88 57.36 64.04
Causal modeling approaches
VCSR (Wei et al. 2023) 53.00 65.41 - - -
MCR+HCRN (Zang et al. 2023) 49.20 66.00 - - -
VLMs/Multimodal foundation approaches
BLIP-2 (Li et al. 2023) 45.00 62.00 23.32 43.44 53.00
VideoLLaMA (Zhang, Li, and Bing 2023) 42.00 31.00 11.73 28.24 33.31
VideoChat2 (Li et al. 2024) 60.00 46.00 15.36 40.45 48.50
GPT4o (OpenAI 2024) 70.00 52.00 48.17 56.72 58.00
VILA 1.5-3B (Lin et al. 2024) 60.23 72.11 62.80 65.05 67.20
DeepSeek-VL2 (Wu et al. 2024) 51.55 57.08 17.12 41.92 51.95
QwenVL 2.5-7B (Bai et al. 2025) 70.75 76.22 31.54 59.50 70.73
Ours 63.95 76.18 67.65 69.26 71.22
Table 2: Performance evaluation with SOTA methods. nxqa: NextQA; cvqa: CausalVidQA; cchqa: CausalChaos! datasets. WtAvg: Weighted average.
Figure 5: Qualitative results. GT AO: Groundtruth Answer Option; BM: Baseline Model; SA: Selected Answer; $C$: Causal Chain. Only a few frames per video are shown. Green and red boxes indicate success and failure cases. In the first example, actor masks come from the CausalVidQA dataset, which includes reference-based QA. Please zoom-in.
Results.
Table 2 shows that traditional VideoQA models lag behind VLMs. Notably, MIST—though not a large model—performed strongly, likely due to its long-term video understanding design and use of CLIP features with advanced spatiotemporal attention. Moreover, our method is simpler yet more effective, significantly outperforming prior causal models like MCR+HCRN (best performing model in (Zang et al. 2023))—closest to our work. Additionally, our model provides interpretable, symbolic causal chains—distinct from latent-variable or counterfactual approaches. Next-generation VLMs like (Lin et al. 2024), trained on diverse multimodal tasks, outperformed even the closed-source GPT-4o on causal VideoQA. Building on (Lin et al. 2024), our framework outperforms all existing models, likely due to: 1) improved visual grounding via causal chain extraction; and 2) explicit reasoning in the CCDA. In contrast, VLMs perform grounding and reasoning jointly, which may cause them to overlook visual evidence and rely on biases like plausibility or correlation. See Figure 3 for qualitative results. Our model takes a structured approach, guiding the causal chain extractor to capture cause-and-effect relationships in the video. We view this as a dense prediction task, where fine-grained supervision enhances video understanding. In contrast, typical VideoQA models use a single-label objective, which can lead to shallow learning—guessing answers without true video understanding or causal reasoning. Our approach avoids this by isolating video comprehension from bias, allowing the CCDA to focus solely on selecting the correct answer from the extracted causal chains (reasoning steps). Causal chains provide concise, focused context. While the CCDA made fewer errors, the Causal Chain Extractor sometimes misinterpreted object roles or actions despite correct identification, leading to flawed chains and reduced CCDA accuracy (see Figure 4). A good strategy for future research might be pursue improvements in: 1) role or relationship modeling; and 2) situation understanding. We further compare the quality of generated causal chains from our model with a representative SOTA VLM-QwenVL2.5 in Table 3—results averaged across all datasets. We found that our model performs significantly better than VLM counterpart. This indicates that causality aspect maybe overlooked in the current multimodal understanding literature (highlighted in (Ma 2025)), a gap that our work fills. We believe models like ours also have the potential to serve as reasoning engines, which should be explored by future work.
Model B1 B2 B3 B4 M R S CCS
QwenVL2.5-3B Oneshot 0.35 0.23 0.16 0.11 0.54 0.36 0.40 0.75
Ours 0.63 0.47 0.36 0.28 0.61 0.50 0.52 0.89
Table 3: Causal chain generation performance results.
Human Studies
Explainable systems benefit two key user groups: researchers/system designers and consumers, each with distinct objectives. We outline their goals and systematically design human subject studies to evaluate how our explainable system compares to black-box models across various criteria. Most SOTA models/VLMs operate as black boxes. We select VILA (Lin et al. 2024) as a representative black-box model due to its strong performance in previous experiments and lack of explicit model explanations. To minimize bias, all the participants of this study were not a part of the project. Following previous study (Parmar et al. 2024), we perform all user studies with six participants—one more than the previous study.
Study I: Explainability
In this study, participants are presented with 50 questions and models’ outputs, including final answers and, when available, intermediate explanations in the form of causal chains. They rate their understanding of the explanation,i.e., causal chains, regardless of the correctness of the final answer. To avoid the ambiguity of arbitrary scales (e.g., 1–5), we use a comparative evaluation: participants choose between: 1) System A—no explanation (BlackBox model); 2) System B—Our Model with causal chain explanations; or 3) No Preference—indicating no added value from the causal chains as explanations. Table 4 shows participants found causal chains useful explanations in over 69% of cases, demonstrating their overall effectiveness. Nonetheless, in $\sim$29% of cases, the explanations were not considered helpful, suggesting opportunities for further refinement.
Study II: Trustworthiness
Model Explainability Trustworthy Human Preferred
Blackbox 01.33 01.33 14.81
Ours 69.33 62.67 85.18
No Preference 29.33 36.00 n/a
Table 4: Human study results (%) along various axes.
In this study, participants are presented with 50 questions and models’ outputs, including final answers and, when available, intermediate explanations in the form of causal chains. The goal is to evaluate users’ trust in the systems based on the explanations and the perceived correctness of the answers, without access to ground-truth labels. Participants select one of the following options: 1) System A—trust the BlackBox model’s answer; 2) System B—trust Our Model’s answer with its causal chain explanation; or 3) No Preference—no clear preference between the two. Table 4 summarizes that in over 62% of cases, participants expressed significantly greater trust in our model. However, trust declined when the model’s predictions conflicted with participants’ expectations. We anticipate that trust will increase as the accuracy of our approach improves.
Study III: Human Preference
In this study, we analyze human preference between the two systems: 1) BlackBox model; or 2) Our Model. To ensure a fair comparison, we consider only successful predictions (i.e., those matching groundtruth answers) from both models. Human participants evaluate 50 samples, selecting their preferred model. Participants are instructed to assume they are using an AI system designed to study human behavior and explain the “why” behind people’s actions in the videos. The results of this study are presented in Table 4. We found that in over 85% of cases, participants strongly preferred a system that provides explanations, like our approach, to support its decisions.
Study IV: Utility in System Debugging
Model VLM LLM Cannot Tell
Blackbox 15.00 15.00 70.00
Ours 48.33 38.33 13.33
Table 5: Utility from system debugging perspective.
Researchers and system designers, responsible for developing AI systems, aim to evaluate and improve model performance through systematic debugging. Identifying limitations is a key first step, where intermediate outputs—alongside final predictions—offer valuable insights. To test this hypothesis, we designed an experiment targeting failure cases of both our explainable and black-box models. Although it assumes knowledge of incorrect predictions, this reflects a realistic scenario, as system designers typically work with labeled data. Researchers are then asked to identify the source of the model’s failure: 1) visual perception, 2) language and reasoning, or 3) Cannot Tell. The third option indicates uncertainty, while choosing the first two suggests a clearer understanding of the system and supports more confident diagnosis. Six researchers with expertise in computer vision, VLMs, and LLMs each analyze 20 failure cases from both systems, identifying fault locations. For our model, they view the video, question, causal chain explanations, and final predictions; for the black-box model, only the video, question, and final predictions are provided. Participants are instructed to use all available information in their assessments. Table 5 shows that the ”Cannot Tell” option was chosen far less often for our model than in case of BlackBox model. Researchers attributed failures to the chain extractor in $\sim$48% of cases and to the chain-driven answerer in $\sim$38%, suggesting that causal chains improve system debuggability.
Out-Of-Domain Chain Generation
Model B1 B2 B3 B4 S R M CCS
Baseline 0.19 0.10 0.06 0.04 0.35 0.29 0.37 0.42
Ours 0.35 0.22 0.13 0.08 0.39 0.37 0.61 0.84
Table 6: ood chain generation performance evaluation.
Figure 6: Qualitative examples of Out-of-Domain generated causal chains (oodgcc). Please zoom-in.
To test the generalizability of CCE, we train it on a cartoon-based dataset (CausalChaos!) and evaluate it on an out-of-distribution real-world dataset (NextQA), making the task intentionally challenging. Quality of generated chains is evaluated in terms of bleu-1–4 (b), meteor (m), rouge (r), spice (s), our causal coherence score ccs, with groundtruth causal chains (Section 4). A zero-shot vila 1.5 serves as the baseline; our extractor also uses vila 1.5 for a fair comparison. Table 6 shows that our model significantly outperforms the baseline in cross-dataset causal chain generation—suggesting that the reasoning patterns that our causal chain extractor learns are robust and transferable, highlighting its potential as a reusable causal reasoning engine. Qualitative results in Fig 6.
6 Conclusion
We shift paradigms and introduce a principled, structured approach for Causal-Why VideoQA task, which decouples video understanding, reasoning and answer generation. These modules use natural language causal chains as intermediate representations. Our approach enforces analyzing entire videos and explicitly reason to determine the cause behind the actions, instead of shortcut-based answering. At the core of our method lies novel integration of SCMs and CoT. Our approach We show how to efficiently construct training data to develop such models. Thorough experimentation demonstrates that our approach improves explainabiity, performance on VideoQA task, and have the potential to serve as reusable causal reasoning engines.
Acknowledgements. This research/project is supported by the National Research Foundation, Singapore, under its NRF Fellowship (Award# NRF-NRFF14-2022-0001). This research is also supported by funding allocation to B.F. by the Agency for Science, Technology and Research (A*STAR) under its SERC Central Research Fund (CRF), as well as its Centre for Frontier AI Research (CFAR). We would also like to thank Dhruv Verma and Elston Tan.

### v2

@@
-ChainReaction! Structured Approach with Causal Chains as Intermediate Representations for Improved & Explainable Causal Video Question Answering
+ChainReaction: Causal Chain-Guided Reasoning for Modular and Explainable Causal-Why Video Question Answering
 Abstract
@@
 These black-box approaches offer limited interpretability and tend to depend on shallow heuristics.
-We propose a novel, modular framework that explicitly decouples causal reasoning from answer generation, introducing natural language causal chains as interpretable intermediate representations.
+We propose a novel, modular paradigm that explicitly decouples causal reasoning from answer generation, introducing natural language causal chains as interpretable intermediate representations.
 Inspired by human cognitive models, these structured cause-effect sequences bridge low-level video content with high-level causal reasoning, enabling transparent and logically coherent inference.
-Our two-stage architecture comprises a Causal Chain Extractor (CCE) that generates causal chains from video-question pairs, and a Causal Chain-Driven Answerer (CCDA) that produces answers grounded in these chains.
-To address the lack of annotated reasoning traces, we introduce a scalable method for generating high-quality causal chains from existing datasets using large language models.
+Our two-stage architecture comprises a Causal Chain Extractor (CCE) that generates causal chains from video-question pairs, and a Causal Chain-Driven Answerer (CCDA) that derives answers grounded in these chains.
+To address the lack of annotated reasoning traces, we introduce a scalable method for generating accurate causal chains from existing datasets.
+We construct human verified causal chains for 46K samples.
 We also propose CauCo, a new evaluation metric for causality-oriented captioning.
@@
 1 Introduction
-Understanding the motivations behind human actions is crucial for developing advanced systems capable of nuanced behavior analysis.
-Human actions are shaped by various factors, including personal experiences, emotions, social contexts, and cultural backgrounds.
-This complexity demands a comprehensive approach to uncovering the underlying causes of actions.
-In this context, Causal Video Question Answering (Causal-Why Video QA) requires models not just to recognize events, but to explain why they occurred—demanding higher-order reasoning that surpasses the descriptive QA paradigm (Li, Niu, and Zhang 2022; Xiao et al.
-2021; Parmar et al.
-2024; Foss et al.
-2025).
-However, existing Causal-Why Video QA models often draw conclusions from incomplete evidence or rely on shallow heuristics, such as matching action verbs or object nouns in vision-language embedding spaces (Parmar et al.
-2024; Rawal et al.
-2024; Wei et al.
-2023).
-These models typically entangle video understanding, reasoning, and answer generation into a single monolithic process, which makes their reasoning opaque and error-prone.
-Moreover, many high-performing VLM models operate as ”black boxes,” offering limited interpretability into their decision-making processes.
-Figure 1: Concept.
+Understanding the motivations behind human actions is crucial for developing advanced systems for nuanced behavior analysis.
+Human actions are shaped by factors such as personal experience, emotion, social context, and culture.
+This complexity requires uncovering underlying causes.
+In this context, Causal Video Question Answering (Causal-Why VideoQA) asks models not only to recognize events but to explain why they occur—demanding higher-order reasoning beyond descriptive QA [causalvidqa, nextqa, causalchaos, foss2025causalvqa].
+Existing Causal-Why VideoQA models often reason from incomplete evidence or rely on shallow heuristics (e.g., matching action verbs or object nouns in vision-language embedding spaces [causalchaos, rawal2024dissecting, wei2023visual]).
+These models entangle video understanding, reasoning, and answer generation into one monolithic process, making their reasoning opaque and error-prone.
+Many high-performing vision language models (VLMs) also operate as black boxes, offering limited interpretability into their decisions.
+Figure 1: (Top) Concept.
 (1) Existing Video ($\mathcal{V}$) Question ($\mathcal{Q}$) Answer ($\mathcal{A}$) approaches through the lens of structural causal models (SCMs), highlighting their monolithic and black-box nature.
-(2) In contrast, we propose a principled departure from this paradigm: a mediation mechanism based on natural language causal chains ($\mathcal{C}$).
+(2) In contrast, we propose a principled departure from this paradigm: leveraging the Causal Reasoning Trace ($\mathcal{C}$), a structured intermediate representation based on natural language causal chains.
 We factorize this SCM into two SCMs (3,4)—enabling structured video understanding, reasoning, and inference—leading to superior explainability and performance.
+(Bottom) Example.
+Please zoom in for the best view.
 In this work, we make the case that reasoning and answering should be explicitly decoupled and modularized.
-We introduce a new paradigm in which causal reasoning and answer generation are handled by separate modules that communicate via natural language causal chains (Hanson 1955; Pearl and Mackenzie 2018)—structured sequences of cause-effect events that serve as intermediate reasoning steps (Figure 1).
-Causal chains provide logically coherent representations that bridge low-level video content with high-level causal understanding (Trabasso and Van Den Broek 1985).
-To our knowledge, ours is the first approach in Causal-Why VideoQA to explicitly use causal chains as interpretable intermediates, inspired by their role in human cognition and scientific explanation (Gopnik et al.
-2004).
+We introduce a new paradigm in which causal reasoning and answer generation are handled by separate modules that communicate via natural language causal chains [hanson1955causal, pearl2018book]—structured sequences of cause-effect events that serve as intermediate reasoning steps (Figure 1).
+Causal chains or causal reasoning traces provide a logically coherent bridge between low-level video content and high-level causal understanding [trabasso1985causal].
+Formulated as natural-language sequences, they capture the linear, observable steps linking a cause (e.g., a character’s intention) to its effect, ensuring that answers are grounded in the video’s causal progression.
+To our knowledge, ours is the first approach in Causal-Why VideoQA to explicitly use causal chains as interpretable intermediates, inspired by their role in human cognition and scientific explanation [gopnik2004theory].
 At the core of our method lies the integration of Structural Causal Models (SCMs) and Chain-of-Thought (CoT) reasoning, which enables a principled, structured decomposition of the VideoQA task.
 This synergy allows us to model causality in a robust and interpretable way, going beyond latent embeddings to explicitly capture causal semantics.
-Our model consists of two stages: 1) a causal chain extractor (Figure 1(3)), and 2) a causal chain-driven answerer (Figure 1(4)).
-The causal chain extractor model learns to extract causal chains from video, conditioned on the causal-why questions.
-The causal chain-driven answerer model learns to generate answers to causal-why questions based on the extracted causal chains.
-However, a key challenge lies in training causal chain extractor.
-In particular, we do not have reasoning traces for training causal chain extractor.
+Our model consists of two stages: 1) a causal chain extractor—CCE—(Figure 1(3)), and 2) a causal chain-driven answerer—CCDA—(Figure 1(4)).
+The CCE model learns to extract causal chains from video, conditioned on the causal-why questions.
+The CCDA model learns to generate answers to causal-why questions based on the extracted causal chains.
+Distinct from general causal inference models that reason over partially observed or hypothetical systems, the CCE operates in fully observed video settings, where the causal process has already unfolded.
+After observing the complete video, the CCE identifies the actualized, linear sequence of events that led to the outcome—making the task descriptive and post-hoc rather than inferential, and emphasizing interpretability and causal fidelity within the VideoQA framework.
+However, a key challenge lies in training CCE.
+In particular, there are no datasets containing reasoning traces for training CCE.
 To tackle this challenge, we develop an approach to generate causal chains from existing VideoQA datasets efficiently.
@@
 Our main contributions can be summarized as:
-  • Proposing a structured framework for Causal-Why Video Question Answering that leverages natural language causal chains as intermediate representations to enhance reasoning and transparency.
+  • Proposing a structured paradigm for Causal-Why Video Question Answering that leverages natural language causal chains as intermediate representations to enhance reasoning and transparency.
   • Introducing a two-stage architecture—Causal Chain Extractor (CCE) and Causal Chain-Driven Answerer (CCDA)—that decouples video understanding from causal inference.
-  • Presenting a novel, scalable and efficient method for generating causal chain annotations using large language models, enabling the augmentation of existing datasets.
+  • Introducing a human-in-the-loop framework that uses large language models to propose causal chain drafts, which are subsequently verified and finalized by human annotators, yielding high-quality causal reasoning data.
   • Introducing CauCo score, a causality-oriented captioning metric.
@@
 2 Related Work
-Chain-of-Thought (CoT)
-is a prompting technique that improves large language model performance by decomposing tasks into intermediate reasoning steps.
-It has proven effective in complex tasks involving arithmetic, logic, and commonsense reasoning, particularly through few-shot prompting with exemplar rationales (Wei et al.
-2022).
-Unlike most CoT-based methods that treat reasoning as latent or emergent, we model reasoning steps—causal chains—as explicit, structured, and learnable variables.
+Chain-of-Thought (CoT) is a prompting technique that improves LLMs by decomposing tasks into intermediate reasoning steps.
+It is effective in arithmetic, logic, and commonsense tasks, especially via few-shot prompting with exemplars [wei2022chain].
+Unlike CoT methods that treat reasoning as emergent, we model reasoning steps—causal chains—as explicit, structured variables.
 This enables supervision, interpretability, and integration with modular architectures, breaking from the monolithic CoT paradigm.
 Causal Reasoning and Structural Causal Models (SCMs).
-Causality has long been a cornerstone of scientific reasoning, formalized in AI through SCMs and do-calculus (Pearl 2009).
-Recent work highlights the importance of causal representation learning for improving robustness and generalization in AI systems (Schölkopf et al.
-2021).
-While SCMs and causal inference methods have been applied in contexts like debiasing (Tang et al.
-2020) and counterfactual explanations (Narendra et al.
-2018), they have rarely been extended to VideoQA.
-We address this gap by integrating SCMs into a modular framework where causal chains derived from video content support interpretable answer generation.
-This positions our approach as a novel instantiation of structured causal reasoning in a domain where temporal complexity often hinders transparency.
+Causality is foundational in scientific reasoning and is formalized in AI through SCMs and do-calculus [pearl2009causality].
+Recent work emphasizes causal representation learning for robustness and generalization [scholkopf2021toward], and SCM-based methods have been used for tasks like debiasing [tang2020unbiased] and counterfactual explanations [narendra2018explaining].
+However, such approaches are rarely applied to VideoQA.
+We address this gap by integrating SCMs into a modular framework where video-derived causal chains support interpretable answer generation—bringing structured causal reasoning to a domain where temporal complexity often obscures transparency.
 Causal Chains and Structured Reasoning in AI.
-Causal chains are central to human explanations, offering intuitive, step-wise accounts of how and why events unfold (Parmar et al.
-2024; Niu et al.
-2025).
-In AI, structured representations like causal graphs have shown promise for improving reasoning and interpretability.
-Prior efforts, such as (Wang et al.
-2024; Ayazoglu et al.
-2013), investigate causal interactions in videos or use interventions for action recognition.
-However, these methods stop short of leveraging causal chains as intermediates for reasoning tasks.
-Our work formalizes causal chains as explicit bridges between raw video observations and high-level causal understanding, enabling a principled and interpretable approach to Causal-Why Video QA.
-Video Causal Reasoning and Temporal Understanding.
-A number of recent efforts explore causality in video understanding.
-For example, (Li et al.
-2020; Yang et al.
-2021) employ structural models to identify causal interactions in dynamic systems and moment retrieval.
-However, these do not address question answering.
-Some methods identify key video segments for question answering (Wei et al.
-2023), while others model object interactions over space and time using generative models (Chen et al.
-2023).
-Yet, causal reasoning remains largely implicit or limited to local interactions.
-Most relevant to our work is (Chen et al.
-2025), who propose a dataset and task for identifying event-level causal diagrams in videos.
-Their model infers why a final event occurred but does not generate full causal chains or use them for causal question answering.
-By contrast, we extract event-level causal chains to directly power a dedicated inference module.
+Causal chains provide intuitive, stepwise explanations of how events unfold [causalchaos, niu2025r].
+In AI, structured representations like causal graphs improve reasoning and interpretability, and prior work has explored causal interactions in videos or intervention-based action recognition [wang2024modeling, ayazoglu2013finding].
+However, these approaches do not use causal chains as intermediates for reasoning.
+We instead formalize causal chains as explicit bridges between video observations and high-level causal understanding, enabling a principled and interpretable framework for Causal-Why Video QA.
+Video Causal Reasoning & Temporal Understanding.
+Recent work explores causality in video understanding—e.g., structural models for causal interactions and moment retrieval [li2020causal, yang2021deconfounded]—but not for QA.
+Other methods select key segments for QA [wei2023visual] or model object interactions generatively [chen2023llcp], yet their causal reasoning remains implicit or localized.
+Chen et al.
+[chen2025mecd] introduce event-level causal diagram annotation, but their models explain a final event rather than generate causal chains or use them for QA.
+In contrast, we extract event-level causal chains and leverage them in a dedicated inference module.
 Zang et al.
-(Zang et al.
-2023) explore video-text causal links, while Su et al.
-(Su et al.
-2023) generate causal questions from captions, but they do not model causal chains or explicitly decouple reasoning from answering.
-Vision-Language Models and Their Limitations in Causality.
-Recent Vision-Language Models (VLMs) such as VideoLLaMA (Zhang, Li, and Bing 2023), VideoChat2 (Li et al.
-2024), and VILA (Lin et al.
-2024) have advanced video QA and captioning by leveraging large-scale video-text pretraining.
-However, these models focus on recognizing and describing events, not explaining them.
-Their monolithic architectures lack interpretability and often rely on shallow correlations.
-Our approach complements these models by introducing causal chains as structured reasoning intermediates—enhancing both performance and transparency in Causal-Why Video QA.
+[zang2023discovering] study video–text causal links, and Su et al.
+[su2023language] generate causal questions from captions, but neither models causal chains or decouples reasoning from answering.
+Vision-Language Models (VLMs) & their Limitations in Causality.
+Recent VLMs (e.g., VideoLLaMA [videollama], VideoChat2 [videochat2], VILA [vila], etc.) advance VideoQA & captioning through largescale video–text pretraining, but primarily recognize and describe events rather than explain them.
+Their monolithic architectures offer limited interpretability & often rely on shallow correlations.
+Our method complements them by introducing causal chains as structured reasoning intermediates, improving both performance & transparency in Causal-Why Video QA.
+3 Causal Chain Construction for SFT
+We propose a novel, scalable and efficient methodology to construct causal chains to be used for supervised finetuning (SFT) of our causal chain extractor (Section 4).
+We generate causal chains that accurately reflect the reasoning behind human-written answers to video-based questions.
+Base datasets: since no existing VideoQA dataset includes causal chain annotations to support our approach, we collect causal chains for three challenging causal video QA datasets: 1) NextQA [nextqa], 2) CausalVidQA [causalvidqa], and 3) CausalChaos!
+[causalchaos].
+We refer to these three datasets as the base (or source) datasets.
+We focus on Causal-Why QA in these base datasets.
+In base datasets, each sample is a paired triplet of {Video ($\mathcal{V}$), Question ($\mathcal{Q}$), Answer ($\mathcal{A}$)}.
+Causality defined in the base datasets: the causal-why questions and correct answers in base datasets are human-authored, as a result the notion of cause or causality is well defined in the data itself.
+All the annotations are rigorously cross-annotator verified, resulting in multihuman agreement on causality.
+Our goal is to learn to reason through these causal relations.
+We have provided further details on these three base datasets in the Appendix.
+Preliminary I.
+Human annotators watch the video ($\mathcal{V}$), formulate a question ($\mathcal{Q}$) about it, and write the correct/gold answer ($\mathcal{A}$).
+The annotations are cross-verified, ensuring multi-annotator agreement.
+Annotators implicitly use causal chains ($\mathcal{C}$) when writing answers, as illustrated in Figure 2(1).
+Thus, the detailed QA pairs of our base datasets contain causal chains embedded within them.
+Preliminary II.
+Given the questions and the corresponding correct, gold answers, LLMs have been demonstrated to accurately recover the intermediate reasoning steps or causal chains.
+Illustrated in Figure 2(2).
+Corollary I.
+Based on preliminaries I and II, we propose that if provided with the questions and the corresponding human written correct, gold answers from the base datasets, then powerful oracle LLMs (e.g., [gpt4o, deepseekvl2, gemini]) can reliably recover the reasoning steps or causal chains.
 Figure 2: Causal chain construction for SFT.
@@
 (2) We propose to recover these causal chains with the help of LLM using questions and correct gold answers.
-(3,4) Our robust causal chain generation and verification pipeline.
-3 Causal Chain Construction for SFT
-In this section, we propose a novel, scalable and efficient methodology for the construction of causal chains to be used for supervised finetuning (SFT) of our causal chain extractor (presented in later sections).
-Our goal is to generate causal chains that accurately reflect the reasoning behind human-written answers to video-based questions.
-Since no existing VideoQA dataset includes causal chain annotations to support our approach, we collect causal chains for three challenging causal video QA datasets: 1) NextQA (Xiao et al.
-2021), 2) CausalVidQA (Li, Niu, and Zhang 2022), and 3) CausalChaos!
-(Parmar et al.
-2024).
-We refer to these three datasets as the base (or source) datasets.
-We focus on Causal-Why QA in these base datasets.
-In base datasets, each sample is a paired triplet of {Video ($\mathcal{V}$), Question ($\mathcal{Q}$), Answer ($\mathcal{A}$)}.
-We have provided summaries of these three base datasets in the Appendix.
-Preliminary I.
-Human annotators of the base dataset watch the video ($\mathcal{V}$), then formulate a question ($\mathcal{Q}$) related to video; and then finally, meticulously write the correct, gold answer ($\mathcal{A}$) to the question.
-Further, the annotations are rigorously cross-annotator verified, resulting in multihuman agreement.
-Human annotators intuitively and implicitly make use of causal chains ($\mathcal{C}$) in the process of writing answers.
-This process is illustrated in Figure 2(1).
-As a result, the detailed QA pairs of our base datasets contain causal chains embedded within them.
-Preliminary II.
-Given the questions and the corresponding correct, gold answers, LLMs have been demonstrated to accurately recover the intermediate reasoning steps or causal chains. Illustrated in Figure 2(2).
-Corollary I.
-Based on preliminaries I and II, we propose that if provided with the questions and the corresponding human written correct, gold answers from the base datasets, then powerful oracle LLMs (e.g., GPT4o, DeepSeek, Gemini) can reliably recover the reasoning steps or causal chains.
-Causal Chain Construction Methodology.
-Following Corollary I, we propose a robust methodology for causal chain generation and verification in the following.
-  1. 1.
-    Causal Chain Generation Process: We leverage a powerful LLM (GPT4o).
+(3,4) Our robust causal chain generation and manual verification and video grounding check pipeline.
+Causal Chain Construction Methodology. Following Corollary I, we propose a robust methodology for causal chain generation and verification in the following.
+  1. wide
+    Causal Chain Draft Generation Process: We leverage a powerful LLM (GPT4o).
 We refer to it as Oracle LLM.
@@
 Illustrated in Figure 2(3).
-  2. 2.
+Oracle LLM is not used as end-task solvers but as structured reasoning components within a rigorously controlled pipeline.
+Their role is analogous to that of annotator assistants—providing drafts that undergo multi-layer human validation to ensure reliability and reproducibility.
+  2. wiide
     Causal Chain Verification Process: Illustrated in Figure 2(4).
 We ensure that the generated causal chains are of high quality and accurately capture the reasoning steps using rigorous Quality Checks consisting of:
-    1. (a)
-      Programmatic Validation: Programs are developed for checking proper structure, formatting, completeness and length (constraint: $\leq$ 10 events) of the generated causal chains.
-    2. (b)
+    1. wide
+      Programmatic Validation: Programs check generated causal chains for structure, format, completeness, and length ($\leq$10 events).
+    2. wiide
       Cross-LLM Verification: Verifier LLM is provided with question ($\mathcal{Q}$), correct gold answer ($\mathcal{A}$) and the generated causal chain ($\mathcal{C}$) Generated causal chains are independently reviewed and verified by second LLM to assess the chain’s logical coherence, relevance, and consistency with $\mathcal{Q}$ and $\mathcal{A}$ pair.
@@
 Specifically, we use Gemini 2.5 as the verifier LLM.
-    3. (c)
-      Manual Verification against $\mathcal{V}$: Human verifiers are provided with full access to the corresponding video $\mathcal{V}$, the question $\mathcal{Q}$, and the correct gold answer $\mathcal{A}$ to ensure proper grounding of the causal chain in the video content.
-Chains passing the first two stages are verified by humans for logical coherence, relevance, and consistency with $\mathcal{Q}$ and $\mathcal{A}$ pair and grounded in video ($\mathcal{V}$).
-Human verifiers are computer science graduates and well-versed with the task and literature.
-Human verifiers are allowed to manually add missing details to the causal chains.
-In addition, human verifiers are trained to identify hallucinations (e.g., events not observable or implied in the video) during manual verification.
-Chains containing any such hallucinations are rejected and regenerated.
-    If the chain fails at any stage, it is regenerated, until it passes all the checks.
+    3. wiiide
+      Manual Correction & Verification against $\mathcal{V}$: Human verifiers receive the video $\mathcal{V}$, question $\mathcal{Q}$, and correct gold answer $\mathcal{A}$ to ensure proper grounding of the causal chain in the video content.
+Chains passing the first two stages are checked for logical coherence, relevance, and consistency with $\mathcal{Q}$, $\mathcal{A}$, and $\mathcal{V}$.
+Verifiers—computer science graduates familiar with the task—may add missing details to causal chains and flag hallucinations (e.g., events not observable or implied in the video).
+Chains with hallucinations or other failures are regenerated until they pass all checks.
+Only the chains that have passed all the checks are considered for the final dataset.
 Finally, 1000 samples belonging to each base dataset are manually reviewed by the authors.
-On average, over 95% of the chains passed the author verification, ensuring high-quality annotations.
-Our methodology allowed to generate consistent and reliable causal chains with respect to the video context.
-In total, we collected causal chains for 46024 samples across three base datasets.
-Further statistics provided in Appendix.
+Over 95% of the chains passed the author verification, ensuring accurate causal chain annotations.
+Our method produced reliable, accurate causal chains grounded in the video context.
+In total, we constructed human-verified causal chains for 46,024 samples across three datasets.
+Further stats provided in the Appendix.
+Dataset will be released for reproducibility and future research.
 4 Approach
-Theoretical Foundation and Motivation
-Our work is grounded in the principles of Structural Causal Models (SCMs) (Pearl 2009) and Chain-of-Thought (CoT) (Wei et al.
-2022) reasoning.
+4.1 Overview and Motivation
+Conceptually, our work is grounded in the principles of Structural Causal Models (SCMs) [pearl2009causality] and Chain-of-Thought (CoT) [wei2022chain] reasoning.
 We decompose the VideoQA task into two explicit stages: 1) causal chain extraction; and 2) causal chain-driven answering.
@@
 From an SCM perspective, we introduce causal chains as structured intermediate representations.
-While many existing VideoQA approaches can be abstracted as an undifferentiated model where video ($\mathcal{V}$) and question ($\mathcal{Q}$) jointly influence the answer ($\mathcal{A}$) as ($\mathcal{V}\rightarrow\mathcal{A}\leftarrow\mathcal{Q}$), we propose a structured mediation via a causal chain variable ($\mathcal{C}$), yielding the following SCM: $\mathcal{V}\rightarrow\mathcal{C}\rightarrow\mathcal{A},\mathcal{Q}\rightarrow\mathcal{C}\rightarrow\mathcal{A}$.
-Then, we factorize this SCM into: $\mathcal{V}\rightarrow\mathcal{C}\leftarrow\mathcal{Q}$ (this becomes our Causal Chain Extractor—Figure 1(3)) and $\mathcal{Q}\rightarrow\mathcal{A}\leftarrow\mathcal{C}$ (this becomes our Causal Chain-driven Answerer—Figure 1(4)).
+While many existing VideoQA approaches can be abstracted as an undifferentiated model where video ($\mathcal{V}$) and question ($\mathcal{Q}$) jointly influence the answer ($\mathcal{A}$) as ($\mathcal{V}\rightarrow\mathcal{A}\leftarrow\mathcal{Q}$), we propose a structured reasoning via a Causal Reasoning Trace ($\mathcal{C}$), yielding the following SCM: $\mathcal{V}\rightarrow\mathcal{C}\rightarrow\mathcal{A},\mathcal{Q}\rightarrow\mathcal{C}\rightarrow\mathcal{A}$.
+Then, we factorize this SCM into: $\mathcal{V}\rightarrow\mathcal{C}\leftarrow\mathcal{Q}$ (this becomes our Causal Chain ExtractorFigure 1(3)) and $\mathcal{Q}\rightarrow\mathcal{A}\leftarrow\mathcal{C}$ (this becomes our Causal Chain-driven AnswererFigure 1(4)).
 Causal Chain Extractor module would first extract $\mathcal{C}$, and pass it to the Causal Chain-Driven Answerer module.
-This structured factorization aligns with CoT principles by making reasoning steps explicit, and reflects SCM modularity by separating causal understanding from answer generation.
+We assume access to fully observed videos, where the causal process leading to an outcome manifests as a temporally ordered sequence of events—a single realized traversal through the underlying causal graph.
+Once this process is complete, the relevant causal information resides entirely within this factual path.
+Thus, for Causal-Why VideoQA, the causal chain can be modeled as a linear sequence of observed cause–effect events, providing a complete & interpretable representation of the reasoning process.
+Detailed treatment of the proposed use of linear causal chains is provided in the Appendix.
+The structured factorization aligns with CoT principles by making reasoning steps explicit, & reflects SCM modularity by separating causal understanding from answering.
 As a result, our approach enables:
-  • Focused processing: Unlike typical VideoQA approaches, the proposed structured approach allows more focused processing at each stage, with the outputs of each stage explicitly made available to the subsequent one.
-By handling smaller, more focused chunks, there is a lesser risk of overlooking important reasoning steps or drawing incorrect conclusions unlike existing approaches.
-This approach draws inspiration from the Chain-of-Thought philosophy, which we adapt into a structured model.
+  • Focused processing: Unlike existing VideoQA methods, our structured approach enables focused processing, with each stage’s output passed to the next.
+Processing smaller chunks reduces the risk of missing reasoning steps or making incorrect inferences.
+It is inspired by the Chain-of-Thought philosophy, adapted into a structured model.
   • Improved video understanding: Modeling cause-and-effect relationships explicitly in causal chains is a dense prediction task.
-With detailed supervision, vision models learn to better able to capture rich video content.
-In contrast, typical VideoQA models are trained with a single label (the correct answer), which may be insufficient to fully characterize the video’s complexity (Bi et al.
-2025).
+With detailed supervision, vision models learn to better capture rich video content.
+In contrast, typical VideoQA models are trained with a single label, which may be insufficient to capture the video’s complexity [bi2025reasoning].
   • Enhanced explainability: By generating human-readable causal chains as intermediate outputs, our model improves both explainability & interpretability, making the reasoning process more transparent.
-Model
-Thus far, we saw that our approach factorizes VideoQA problem into subproblems: causal chain extraction and causal chain-driven answerer.
+4.2 Model
+Thus far, we saw that our approach factorizes VideoQA problem into modules: causal chain extraction and causal chain-driven answerer.
 Now, we explain these individual modules and their training/inference in the following.
-Causal Chain Extractor (CCE)
+4.2.1 Causal Chain Extractor (CCE)
 The objective of CCE module is to explicitly capture the cause-and-effect relationships or the reasoning steps from videos, conditioned on the questions, and express them in a detailed yet concise form.
-Generating causal chains from videos conditioned on questions constitutes a complex reasoning task.
-Towards that end, we propose to leverage the representational capacity of foundation models.
-However, generating causal chains is non-trivial task that extends beyond the native capabilities of existing vision-language foundation models.
-Although foundation models exhibit strong performance across a wide range of vision-language benchmarks, they demonstrate limited proficiency in causal chain generation, primarily due to the absence of structured causal supervision in their pretraining corpora.
-To address this limitation, a novel dataset comprising video–question–causal chain triplets is introduced, as described in Section 3.
-This dataset enables the supervised fine-tuning (SFT) of pretrained vision-language foundation models to align their generative capacity with the demands of structured causal inference.
-Through this process, the model is explicitly trained to capture event-level dependencies and produce human-interpretable causal chains grounded in both visual and linguistic context.
+The CCE extracts the causal chain after the entire video has been observed, when the full causal process and its outcome are known.
+At this stage, the causal structure is realized as a single factual sequence of events leading to the answer.
+The extractor therefore identifies this linear causal path directly, without reasoning over counterfactual or hypothetical alternatives.
+This formulation treats causal chain extraction as a descriptive recognition task, emphasizing interpretability, fidelity, and robustness within the VideoQA paradigm.
+Generating causal chains from videos conditioned on questions is a complex reasoning task.
+We leverage the representational capacity of foundation models.
+However, generating causal chains is non-trivial and beyond existing vision-language foundation models.
+Although foundation models exhibit strong performance across vision-language benchmarks, they lack structured causal supervision, limiting causal chain generation.
+To address this limitation, we introduce a novel dataset of video–question–causal chain triplets (Section 3).
+This dataset enables supervised fine-tuning (SFT) of pretrained vision-language models for structured causal inference.
+This trains the model to capture event-level dependencies and generate human-interpretable causal chains grounded in visual and linguistic context.
 Formally, given a video $\mathcal{V}$ and a corresponding question $\mathcal{Q}$, the model learns a mapping: $f_{CCE}:(\mathcal{V},\mathcal{Q})\rightarrow\mathcal{C}$; where $\mathcal{C}$ denotes the generated causal chain and $f_{CCE}$ represents the fine-tuned vision-language foundation model.
-Figure 3: Training of both modules in stages.
-Causal Chain-Driven Answerer (CCDA)
+Further model implementation details provided in Appendix.
+Figure 3: Stage-wise training of our model.
+4.2.2 Causal Chain-Driven Answerer (CCDA)
 The objective of CCDA module is to select the correct answer from candidate answers, based on the questions and the causal chains.
 Existing large language models have shown strong performance on a variety of language-based tasks.
-We propose leveraging language models to implement the Causal Chain-Driven Answerer (CCDA).
+We propose leveraging language models to implement the CCDA.
 However, processing causal chains using language models is non-trivial and mostly not covered in their training suites, largely due to lack of explicit causal chains or step-by-step annotations of causal reasoning.
@@
 Notably, the Causal Chain Extractor (CCE) remains frozen at this stage.
-Formally, given a question $\mathcal{Q}$, the corresponding causal chain $\mathcal{C}$, and answer options $\mathcal{O}$ the model learns a mapping: $f_{CCDA}:(\mathcal{Q},\mathcal{C},\mathcal{O})\rightarrow A$; where $\mathcal{A}$ denotes the selected answer option and $f_{CCDA}$ represents a finetuned language model.
-Inference
+Formally, given a question $\mathcal{Q}$, the corresponding causal chain $\mathcal{C}$, and answer options $\mathcal{O}$ the model learns a function: $f_{CCDA}:(\mathcal{Q},\mathcal{C},\mathcal{O})\rightarrow\mathcal{A}$; where $\mathcal{A}$ denotes the selected answer option and $f_{CCDA}$ represents a finetuned language model.
+4.2.3 Clean Separation with Stage-wise Training
+Problem with end-to-end (E2E) training.
+E2E VQA models optimize all components—from feature extraction to answer prediction—using final answer loss.
+This causes gradient leakage, letting error signals flow through the entire pipeline.
+As a result, models may learn “shortcuts” that boost final answer accuracy but harm causal reasoning.
+Our solution—stagewise training. Our two-stage, stagewise training prevents this by introducing a clear boundary:
+  • CCE Training: The CCE is trained independently to match the ground-truth causal chains ($C$), using a loss defined on $C$ that enforces accurate causal grounding.
+  • CCDA Training: The CCDA is trained on the CCE’s output, with the CCE’s weights typically frozen or its predictions treated as fixed during this stage.
+By freezing the CCE, no gradients can pass backward from the CCDA’s answering loss to the CCE’s internal weights.
+This ensures that the CCE’s learned causal reasoning logic remains pure and untainted by the pressure to maximize the answer score, thus preserving the integrity, robustness, and causal semantics of the intermediate representation.
+4.2.4 Inference
 During inference, the CCE predicts a causal chain from the video and question without using ground-truth chains.
 The CCDA then uses this chain, the question, and answer options to select the correct answer.
-Causality-aware Captioning Metric
-To evaluate the quality of generated causal chains, we also assess how respectful the chains are in terms of modeling causality.
-Since regular captioning metrics do not capture this aspect, we introduce a novel Causal Coherence Score—CauCo score, as an indicator of whether causal chains are causally coherent, i.e., events in the chains are logically connected by cause and effect.
-To achieve this, following a similar approach of LLM as verifiers, we SFT a LLM to identify if a chain is causally coherent.
-The final expected output of this evaluator is expected to be ”True” or ”False”.
-For SFT, we construct a set consisting of positive and negative samples.
-The positive samples are correct causal chains.
-To generate negative samples, we perturb the correct chains with a possibility of six distinct strategies, namely: actor swapping, event negation, event removal, event order reversal, semantic modification, and chain shuffling.
+4.3 CauCo: A Causal Coherence Metric
+We introduce the CauCo score to evaluate how well generated causal chains model causality—an aspect regular captioning metrics miss.
+CauCo quantifies causal consistency and provides causal guarantees in open-world settings by measuring whether events are logically linked by cause and effect.
+Following the LLM-as-verifier approach, we SFT an LLM to judge whether a chain is causally coherent.
+The evaluator outputs “True” or “False.” For SFT, we construct positive and negative samples: positive ones are correct causal chains, and negatives are created by perturbing them using six strategies—actor swapping, event negation, event removal, event order reversal, semantic modification, and chain shuffling.
 5 Experiments
@@
 Implementation details.
-We use PyTorch (Paszke et al.
-2019) to implement models.
+We use PyTorch [paszke2019pytorch] to implement models.
 Noting the strong performance of VILA 1.5 and LLaMA on vision and language tasks, we adopt them as representative models in implementing our CCE and CCDA, respectively.
-Note that our approach is not designed for or limited to any models; practitioners may use models of their choice such as (Zhang, Li, and Bing 2023; Li et al.
-2024; Lin et al.
-2024; Chen et al.
-2024).
-Our CCE is based on VILA-3B (Lin et al.
-2024); and our CCDA is LLaMA-3.1-8B (Grattafiori et al.
-2024) version.
+Note that our approach is not designed for or limited to any models; practitioners may use models of their choice such as [videollama, videochat2, vila, internvl25].
+Our CCE is based on VILA-3B [vila]; CCDA is LLaMA-3.1-8B [grattafiori2024llama] version.
 Further implementation details are provided in Appendix.
-We will release the codebase along with causal chains.
-Datasets. Unless specified otherwise, we conduct experiments on the datasets as discussed in Section 3.
+Codebase along with causal chains will be released.
+Datasets. We conduct experiments on the datasets as discussed in Section 3.
 Tasks. We evaluate on a multiple-choice QA task with five answer options per question, only one of which is correct.
 Performance metric.
-Following prior work (Xiao et al.
-2021; Li, Niu, and Zhang 2022; Parmar et al.
-2024), we use Accuracy as the performance metric.
-For causal chain quality, we use our CauCo score and captioning metrics: BLEU (Papineni et al.
-2002), Meteor (Banerjee and Lavie 2005), ROUGE (Lin 2004), SPICE (Anderson et al.
-2016) scores.
-Experimental Upper bound on Performance
-NextQA CVQA CausalChaos! Avg
-99.70 99.85 98.65 99.40
+Following prior work [nextqa, causalvidqa, causalchaos], we use Accuracy as the performance metric.
+For causal chain quality, we use our CauCo score and captioning metrics: BLEU [bleu], Meteor [meteor], ROUGE [rouge], SPICE [spice].
+5.1 Experimental Upper bound on Performance
+Dataset NextQA CVQA CausalChaos! Overall
+CCDA Accuracy 99.70 99.85 98.65 99.40
 Table 1: Experimental results (Accuracy in %) for answering based on ground truth causal chains.
-Figure 4: Sensitivity to causal chains.
-(1) plot of performance drop vs links removed.
-(2) Qualitative example.
-OCC: original causal chains, MCC: masked chains, SA: selected answer.
-(3) Qualitative examples of SA changes based on hypothetical chains (HCC).
-Please zoom-in.
 We first investigate whether causal chains can serve as effective intermediate representations.
@@
 This finding suggests that using causal chains as intermediate representations is a promising paradigm worth further exploration.
-Probing sensitivity to causal chains.
-We investigated how the CCDA depended on causal chains; and how sensitive it is to the quality of causal chains.
-For this, we incrementally mask-out links of the causal chains.
-Quantitative and qualitative results in Figure 4(1,2) show how the performance degrades as masking is increased.
-We also tested how hypothetical causal chains influence CCDA’s answer selection in an intuitive way (Figure 4(3)).
-Performance Comparison with SOTA
+Figure 4: Causal chain ablation study.
+(1) Study-I: QA accuracy drops by 73% when chains are perturbed.
+(2) Study-II: drop in QA is correlated to amount of perturbation.
+(3,4) Qualitative example.
+OCC: original causal chains, MCC: masked chains, SA: selected answer, PCC: perturbed chains.
+SA changes intuitively as chains are perturbed.
+Please zoom-in.
+5.2 Ablation Studies on the Role of Causal Chains
+To test whether CCDA’s improvements arise from genuine causal reasoning rather than surface-level context enrichment, we conducted ablation studies that systematically degrade the quality of causal chains.
+Study-I: We semantically perturbed the causal chains such that contextual information was preserved but causal relations were disrupted.
+We observed a sharp decline in QA accuracy once causal information was perturbed, despite the context remaining intact (Figure 4(1)).
+Study-II: We progressively masked links and measured QA performance; accuracy degraded monotonically with increased perturbation (Figure 4(2)).
+Causal-chain quality strongly correlated with CCDA accuracy (r=0.97).
+These results empirically validate that CCDA’s reasoning depends on causal-chain integrity, with improvements driven by causal reasoning rather than spurious correlations.
+Qualitative examples (Figure 4(3,4)) show how chain perturbations intuitively alter CCDA’s answers.
+5.3 Performance Comparison with SOTA
 Baselines.
-Following prior work (Xiao et al.
-2021; Li, Niu, and Zhang 2022; Parmar et al.
-2024), we compare against a wide range of models: 1) traditional approaches {BlindQA, EVQA, CoMem, HME, HCRN, HGA}; 2) causal approaches {MCR+HCRN, VCSR}; and 3) SOTA VLMs/Multimodal foundation approaches {MIST, BLIP-2, Video-LLaMA, VideoChat2, GPT-4o, VILA, DeepSeekVL2, Qwen-VL2.5} which excel on vision-language tasks.
-Model NXQA CVQA CChQA Avg WtAvg
+Following prior work [nextqa, causalvidqa, causalchaos], we compare against a wide range of models: 1) traditional approaches [evqa, comem, hme, hcrn, hga]; 2) causal approaches [wei2023visual, zang2023discovering]; and 3) SOTA VLMs/Multimodal foundation approaches [mist, blip2, videollama, videochat2, gpt4o, vila, deepseekvl2, Qwen2.5-VL], which excel on vision-language tasks.
+Model NeXT-QA [nextqa] CausalVidQA [causalvidqa] CausalChaos!QA [causalchaos] Avg WtAvg
 Traditional VideoQA approaches
-BlindQA (Antol et al. 2015) 28.38 59.46 13.07 33.64 44.87
-EVQA (Antol et al. 2015) 42.31 60.95 13.48 38.91 50.62
-CoMem (Gao et al. 2018) 46.15 62.79 13.88 40.94 53.05
-HME (Fan et al. 2019) 46.52 61.45 14.02 40.66 52.43
-HCRN (Le et al. 2020) 47.00 61.61 17.00 41.86 52.93
-HGA (Jiang and Han 2020) 47.00 63.51 15.36 42.00 53.88
+BlindQA [evqa] 28.38 59.46 13.07 33.64 44.87
+EVQA [evqa] 42.31 60.95 13.48 38.91 50.62
+CoMem [comem] 46.15 62.79 13.88 40.94 53.05
+HME [hme] 46.52 61.45 14.02 40.66 52.43
+HCRN [hcrn] 47.00 61.61 17.00 41.86 52.93
+HGA [hga] 47.00 63.51 15.36 42.00 53.88
 Semi-traditional approaches
-MIST (Gao et al. 2023) 54.79 72.41 44.88 57.36 64.04
+MIST [mist] 54.79 72.41 44.88 57.36 64.04
 Causal modeling approaches
-VCSR (Wei et al. 2023) 53.00 65.41 - - -
-MCR+HCRN (Zang et al. 2023) 49.20 66.00 - - -
+VCSR [wei2023visual] 53.00 65.41 - - -
+MCR+HCRN [zang2023discovering] 49.20 66.00 - - -
 VLMs/Multimodal foundation approaches
-BLIP-2 (Li et al. 2023) 45.00 62.00 23.32 43.44 53.00
-VideoLLaMA (Zhang, Li, and Bing 2023) 42.00 31.00 11.73 28.24 33.31
-VideoChat2 (Li et al. 2024) 60.00 46.00 15.36 40.45 48.50
-GPT4o (OpenAI 2024) 70.00 52.00 48.17 56.72 58.00
-VILA 1.5-3B (Lin et al. 2024) 60.23 72.11 62.80 65.05 67.20
-DeepSeek-VL2 (Wu et al. 2024) 51.55 57.08 17.12 41.92 51.95
-QwenVL 2.5-7B (Bai et al. 2025) 70.75 76.22 31.54 59.50 70.73
+BLIP-2 [blip2] 45.00 62.00 23.32 43.44 53.00
+VideoLLaMA [videollama] 42.00 31.00 11.73 28.24 33.31
+VideoChat2 [videochat2] 60.00 46.00 15.36 40.45 48.50
+GPT4o [gpt4o] 70.00 52.00 48.17 56.72 58.00
+VILA 1.5-3B [vila] 60.23 72.11 62.80 65.05 67.20
+DeepSeek-VL2 [deepseekvl2] 51.55 57.08 17.12 41.92 51.95
+QwenVL 2.5-7B [Qwen2.5-VL] 70.75 76.22 31.54 59.50 70.73
 Ours 63.95 76.18 67.65 69.26 71.22
-Table 2: Performance evaluation with SOTA methods. nxqa: NextQA; cvqa: CausalVidQA; cchqa: CausalChaos! datasets. WtAvg: Weighted average.
+Table 2: Performance evaluation with SOTA methods. WtAvg: Weighted average.
 Figure 5: Qualitative results.
@@
 Results.
-Table 2 shows that traditional VideoQA models lag behind VLMs.
-Notably, MIST—though not a large model—performed strongly, likely due to its long-term video understanding design and use of CLIP features with advanced spatiotemporal attention.
-Moreover, our method is simpler yet more effective, significantly outperforming prior causal models like MCR+HCRN (best performing model in (Zang et al.
-2023))—closest to our work.
-Additionally, our model provides interpretable, symbolic causal chains—distinct from latent-variable or counterfactual approaches.
-Next-generation VLMs like (Lin et al.
-2024), trained on diverse multimodal tasks, outperformed even the closed-source GPT-4o on causal VideoQA.
-Building on (Lin et al.
-2024), our framework outperforms all existing models, likely due to: 1) improved visual grounding via causal chain extraction; and 2) explicit reasoning in the CCDA.
-In contrast, VLMs perform grounding and reasoning jointly, which may cause them to overlook visual evidence and rely on biases like plausibility or correlation.
-See Figure 3 for qualitative results.
-Our model takes a structured approach, guiding the causal chain extractor to capture cause-and-effect relationships in the video.
-We view this as a dense prediction task, where fine-grained supervision enhances video understanding.
-In contrast, typical VideoQA models use a single-label objective, which can lead to shallow learning—guessing answers without true video understanding or causal reasoning.
-Our approach avoids this by isolating video comprehension from bias, allowing the CCDA to focus solely on selecting the correct answer from the extracted causal chains (reasoning steps).
-Causal chains provide concise, focused context.
-While the CCDA made fewer errors, the Causal Chain Extractor sometimes misinterpreted object roles or actions despite correct identification, leading to flawed chains and reduced CCDA accuracy (see Figure 4).
-A good strategy for future research might be pursue improvements in: 1) role or relationship modeling; and 2) situation understanding.
-We further compare the quality of generated causal chains from our model with a representative SOTA VLM-QwenVL2.5 in Table 3—results averaged across all datasets.
-We found that our model performs significantly better than VLM counterpart.
-This indicates that causality aspect maybe overlooked in the current multimodal understanding literature (highlighted in (Ma 2025)), a gap that our work fills.
+Table 2 reveals that traditional VideoQA models lag behind VLMs.
+MIST, a smaller model, performs well—likely from its long-term video understanding and use of CLIP features with spatiotemporal attention.
+Our method further outperforms prior causal models such as MCR+HCRN [zang2023discovering], the closest to our work, while remaining simpler and producing interpretable causal chains, unlike latent-variable or counterfactual methods.
+Next-generation VLMs such as VILA [vila] outperform GPT-4o on causal VideoQA.
+Building on VILA, our framework surpasses prior models through improved visual grounding via causal chain extraction and explicit CCDA reasoning.
+In contrast, VLMs perform grounding and reasoning jointly, often overlooking visual evidence and relying on plausibility biases.
+Our structured model guides the extractor to capture cause–effect relations in videos—a dense prediction task enhanced by fine-grained supervision.
+Typical VideoQA models use a single-label objective, leading to shallow learning, whereas our approach isolates comprehension from bias, allowing the CCDA to focus on selecting answers from extracted causal chains.
+The CCDA made fewer errors, though the extractor occasionally misinterpreted object roles or actions, producing flawed chains and reducing accuracy (Figure 5).
+Based on our analysis, future work should focus on 1) role or relationship modeling and 2) situation understanding.
+We also compare our model’s causal chains with those from QwenVL2.5, a representative SOTA VLM (Table 3, averaged across datasets).
+Our model performs significantly better, suggesting that causality is often overlooked in multimodal understanding [ma2025causal]—a gap our work addresses.
 We believe models like ours also have the potential to serve as reasoning engines, which should be explored by future work.
@@
 Table 3: Causal chain generation performance results.
-Human Studies
-Explainable systems benefit two key user groups: researchers/system designers and consumers, each with distinct objectives.
-We outline their goals and systematically design human subject studies to evaluate how our explainable system compares to black-box models across various criteria.
-Most SOTA models/VLMs operate as black boxes.
-We select VILA (Lin et al.
-2024) as a representative black-box model due to its strong performance in previous experiments and lack of explicit model explanations.
-To minimize bias, all the participants of this study were not a part of the project.
-Following previous study (Parmar et al.
-2024), we perform all user studies with six participants—one more than the previous study.
-Study I: Explainability
+5.4 Human Studies
+Explainable systems benefit two user groups: researchers/ system designers & consumers.
+We design four studies to evaluate our explainable system against black-box models across multiple criteria.
+Since most SOTA VLMs are black boxes, we use VILA [vila] as a representative model for its strong performance & lack of explicit explanations.
+To ensure reliability, we conducted user studies with six participants, exceeding the sample size used in prior work [causalchaos].
+To minimize bias: 1) no study participants were involved in the project; 2) Sample order randomization & blinding (independent, anonymized tests) was used for all participants.
+Further details provided in Appendix.
+5.4.1 Study I: Explainability
 In this study, participants are presented with 50 questions and models’ outputs, including final answers and, when available, intermediate explanations in the form of causal chains.
@@
 Nonetheless, in $\sim$29% of cases, the explanations were not considered helpful, suggesting opportunities for further refinement.
-Study II: Trustworthiness
+5.4.2 Study II: Trustworthiness
 Model Explainability Trustworthy Human Preferred
@@
 We anticipate that trust will increase as the accuracy of our approach improves.
-Study III: Human Preference
+5.4.3 Study III: Human Preference
 In this study, we analyze human preference between the two systems: 1) BlackBox model; or 2) Our Model.
@@
 We found that in over 85% of cases, participants strongly preferred a system that provides explanations, like our approach, to support its decisions.
-Study IV: Utility in System Debugging
+5.4.4 Study IV: Utility in System Debugging
 Model VLM LLM Cannot Tell
@@
 Table 5: Utility from system debugging perspective.
-Researchers and system designers, responsible for developing AI systems, aim to evaluate and improve model performance through systematic debugging.
-Identifying limitations is a key first step, where intermediate outputs—alongside final predictions—offer valuable insights.
-To test this hypothesis, we designed an experiment targeting failure cases of both our explainable and black-box models.
-Although it assumes knowledge of incorrect predictions, this reflects a realistic scenario, as system designers typically work with labeled data.
+Researchers and system designers seek to improve models through systematic debugging.
+Identifying limitations is the first step; intermediate outputs such as causal chains alongside final predictions provide useful insights.
+We tested this by examining failure cases of our explainable and black-box models.
+While this assumes known errors, it reflects real settings where designers work with labeled data.
 Researchers are then asked to identify the source of the model’s failure: 1) visual perception, 2) language and reasoning, or 3) Cannot Tell.
@@
 Researchers attributed failures to the chain extractor in $\sim$48% of cases and to the chain-driven answerer in $\sim$38%, suggesting that causal chains improve system debuggability.
-Out-Of-Domain Chain Generation
+5.5 Out-Of-Domain Chain Generation
 Model B1 B2 B3 B4 S R M CCS
@@
 To test the generalizability of CCE, we train it on a cartoon-based dataset (CausalChaos!) and evaluate it on an out-of-distribution real-world dataset (NextQA), making the task intentionally challenging.
-Quality of generated chains is evaluated in terms of bleu-1–4 (b), meteor (m), rouge (r), spice (s), our causal coherence score ccs, with groundtruth causal chains (Section 4).
+Quality of generated chains is evaluated in terms of bleu-1–4 (b), meteor (m), rouge (r), spice (s), our causal coherence score ccs, w.r.t.
+groundtruth causal chains (Section 3).
 A zero-shot vila 1.5 serves as the baseline; our extractor also uses vila 1.5 for a fair comparison.
 Table 6 shows that our model significantly outperforms the baseline in cross-dataset causal chain generation—suggesting that the reasoning patterns that our causal chain extractor learns are robust and transferable, highlighting its potential as a reusable causal reasoning engine.
-Qualitative results in Fig 6.
+Qualitative results in Figure 6.
 6 Conclusion